# Benchmarking Attacks on Learning with Errors

### v1

@@ Abstract
Lattice cryptography schemes based on the learning with errors (LWE) hardness assumption have been standardized by NIST for use as post-quantum cryptosystems, and by HomomorphicEncryption.org for encrypted compute on sensitive data. Thus, understanding their concrete security is critical. Most work on LWE security focuses on theoretical estimates of attack performance, which is important but may overlook attack nuances arising in real-world implementations. The sole existing concrete benchmarking effort, the Darmstadt Lattice Challenge, does not include benchmarks relevant to the standardized LWE parameter choices—such as small secret and small error distributions, and Ring-LWE (RLWE) and Module-LWE (MLWE) variants. To improve our understanding of concrete LWE security, we provide the first benchmarks for LWE secret recovery on standardized parameters, for small and low-weight (sparse) secrets. We evaluate four LWE attacks in these settings to serve as a baseline: the Search-LWE attacks uSVP [9], SALSA [50], and Cool&Cruel [43], and the Decision-LWE attack: Dual Hybrid Meet-in-the-Middle (MitM) [22]. We extend the SALSA and Cool&Cruel attacks in significant ways, and implement and scale up MitM attacks for the first time. For example, we recover hamming weight $9-11$ binomial secrets for KYBER ($\kappa=2$) parameters in $28-36$ hours with SALSA and Cool&Cruel, while we find that MitM can solve Decision-LWE instances for hamming weights up to $4$ in under an hour for Kyber parameters, while uSVP attacks do not recover any secrets after running for more than $1100$ hours. We also compare concrete performance against theoretical estimates. Finally, we open source the code to enable future research.
@@ 1 Introduction
Cryptosystems based on Learning with Errors (LWE) have emerged as leading contenders. The LWE problem is: given many instances of a random vector $a$ of dimension $n$ along with a noisy inner product of $a$ with a secret vector $s$ modulo $q$, recover $s$. The hardness of LWE depends on the dimension $n$, the modulus $q$, and the distributions that the secret and the error are drawn from.
Several LWE-based cryptosystems, such as CRYSTALS-KYBER [11] were standardized by NIST, the US National Institutes of Standards and Technology [19] in 2022 for industry use in post-quantum cryptography. Furthermore, all publicly available fully homomorphic encryption (HE) libraries rely on the hardness of LWE for their security. HE was standardized by an industry consortium in 2018 [2], with recent updated analysis given in [14]. Since standardization, LWE-based cryptosystems have been incorporated into applications such as the encrypted messaging protocol, Signal [33], and LWE-based HE schemes have been used in production by Microsoft in a password-breach detection scheme [34], and in the finance industry for encrypted compute on sensitive market data [31].
@@
Our Contributions. We provide concrete benchmarks for LWE attacks with parameter choices from two key real-world use cases of LWE: CRYSTALS-KYBER and Homomorphic Encryption. We implement and evaluate four concrete LWE attacks—uSVP, SALSA, Cool&Cruel, and Dual Hybrid MiTM—on these parameter settings. These are the first ever successful secret recovery attacks on standardized LWE parameters (see Table IX): $n=256$, $\log q=12$, Module-LWE of ranks $2$ and $3$ with binomial secret and error distributions for KYBER; and $n=1024$, $\log q=26,29$ with ternary secrets and narrow Gaussian error for HE.
We start with sparse, or low weight, secrets, to benchmark attacks which succeed on these. The challenge is to successfully recover higher Hamming weight secrets. Prior work has demonstrated secret recovery for non-standard parameters, such as dimension $n=100-200$ and large $\log q$, with the goal of decreasing $q$ to make the LWE problem harder. Our proposed benchmark instead fixes $n$ and $q$ as the choices standardized for KYBER and HE, and identifies Hamming weight (or secret sparsity) as the sliding hardness parameter, to be increased to reach general secrets. For example, Table IX shows secret recovery of hamming weight $9-11$ binomial secrets for KYBER ($\kappa=2$) parameters in $28-36$ hours with SALSA and Cool&Cruel, while we find that MitM can solve Decision-LWE instances for hamming weights up to $4$ in under an hour for KYBER parameters, while uSVP attacks do not recover secrets after running for more than $1100$ hours.
In the process of implementing and evaluating attacks, we made many new technical contributions, including:
@@ 2.1 LWE Settings
  • Binomial, ${{\mathfrak{B}}^{\eta}}$: from a centered binomial distribution by sampling $(a_{1}\ldots a_{\eta},b_{1}\ldots b_{\eta})\leftarrow\{0,1\}^{2\eta}$, then outputting $\Sigma_{i=1}^{\eta}(a_{i}-b_{i})$ [11].
@@ 2.2 LWE in the real world
Homomorphic Encryption. Most publicly available HE libraries implement RLWE-based systems and often use small (binary, ternary, narrow Gaussian), sparse secrets and small error. Small, sparse secrets enable bootstrapping for evaluating deep circuits required for deep neural nets, such as in encrypted AI model inference [29]. In some proposed schemes, secrets have only $h=64$ active bits for dimension $n=2^{14}$ [23]. Table 4.2 of [14] gives the highest modulus $q$ that can be safely used for a given $n$ with ternary or narrow Gaussian secret distribution $\chi_{s}=\mathcal{N}(0,3.19)$ and narrow Gaussian error $\chi_{e}=\mathcal{N}(0,3.19)$. These are listed in Table III.
@@ 3.1 Benchmark Settings
We propose a set of practical benchmark settings to encourage concrete evaluations of LWE attack performance, using the following criteria. First, parameters should come from real-world choices, such as narrow secret and error distributions and LWE variants like Ring- and Module-LWE, that are standardized or used in deployed LWE cryptosystems. Prior work shows that such parameter choices—including binary secrets, small error, and use of the ring-LWE variant—may weaken LWE, necessitating further experimental study [12, 28, 17, 47, 43].
The challenges should also have tunable hardness settings. The hardness of LWE problems depends on dimension $n$, modulus size $q$, error distribution $\chi_{e}$, and secret distribution $\chi_{s}$. Generally, larger dimension $n$, smaller modulus $q$, and secret/error distributions with higher standard deviation ($\sigma_{e}$, $\sigma_{s}$) make LWE more difficult. A good set of challenges should fix some parameters and allow others to vary so that concrete attacks can be benchmarked to provide interesting insights.
Using these criteria, we propose two sets of benchmark settings for measuring concrete LWE attack performance based on two important real-world applications of LWE: KYBER and Homomorphic Encryption. Security guidelines have been proposed for each of these, outlining the required LWE distribution, $n$, $q$, $\chi_{s}$, and $\chi_{e}$ for real-world implementations, as discussed in §2.2 and Tables II and III. Since $n$, $q$, $\chi_{s}$, and $\chi_{e}$ are fixed for KYBER and for $n=1024$, $\chi_{s}$ and $\chi_{e}$ are fixed for HE, our challenges vary problem difficulty by changing the number of nonzero secret coordinates, $h$, and modulus, $q$. We divide our proposed benchmark settings into parameters focused on KYBER, using MLWE, and parameters focused on HE, using RLWE, and present the proposed settings in Tables VI and VI. The challenge is to solve LWE problems for larger $h$ in these standard parameter settings.
@@ 3.2 Choosing Attacks to Evaluate
To choose which attacks to evaluate, we focused on: (1) attacks which have open-source implementations, (2) attacks which have already been evaluated in nontrivial settings (e.g. dimension $n>256$), (3) attacks that require less than $750$ GB of RAM (our machine capacity). We chose to evaluate the 4 attacks discussed in Section 4: uSVP, ML, and Cool&Cruel for Search-LWE and Dual Hybrid MitM for Decision-LWE. Our attack evaluation provides a first set of benchmarks in the proposed settings, and future work should improve these and evaluate additional attacks.
@@ 4 Attack Implementations and Innovations
Here, we present details of the attacks we evaluate, as well as the innovations we introduce to make these attacks run on the proposed benchmark settings. Refer to the original attack papers for details. All attacks are implemented in a codebase available at https://github.com/facebookresearch/LWE-benchmarking. Implementations are written in Python and leverage fplll and flatter [48] libraries for lattice reduction (with enumeration as the SVP oracle in BKZ). Since all attacks rely on lattice reduction, benchmarking them using the same lattice reduction implementation allows for fair comparison. Any improvement to lattice reduction would benefit all attacks.
@@
The space spanned by these rows contains an unusually short vector $\begin{pmatrix}s&-e&-1\end{pmatrix}$. Thus lattice reduction recovers the secret once it finds the shortest vector in the lattice.
Implementation details. Our implementation of uSVP multiplies the $I_{n}$ matrix by a factor $\omega$, which balances $s$ and $-e$ in the discovered short vector. $\omega$ is determined by formulae given in [17]. We run lattice reduction using BKZ2.0 [21] and incorporate the improvements suggested in [37, Appendix A.7, p. 17]. Future improvements to our implementation of the uSVP attack could incorporate more advanced BKZ schemes, such as Pump and Jump [51].
@@ 4.2 ML Attack
However, we observe experimentally that since $\delta\bf{e_{i}}$ falls outside of $f$’s training distribution, $f$ does not produce helpful predictions on this input. Thus, we propose an alternative secret recovery method that embodies this concept, while remaining within the data distribution. We call it a “slope distinguisher.” This distinguisher calculates the slopes, or approximations of the derivatives, of model outputs using the formula $\frac{\partial f}{\partial x_{i}}(a)\approx\frac{f(a+\delta e_{i})-f(a)}{\delta}=:\hat{s}_{i}$ for some $a$ drawn from the data distribution. To account for the noisy predictions, we compute many samples of $\hat{s}_{i}$ and take the most frequently appearing value, rounded to the nearest integer. With this, the ML method can recover general secrets. This new distinguisher is illustrated with an example in Figure 1.
@@ 4.4 Dual Hybrid MiTM
ML Best $h$ 9 18 16 8 10 17
Preproc. hrs $\cdot$ CPUs $28$cdot $3216$ $11$cdot $3010$ $33$cdot $1843$ $21.5$cdot $1160$ $31.6$cdot $1164$ $23.8$cdot $1284$
Recover hrs $\cdot$ GPUs $8$cdot $256$ $16$cdot $256$ $6$cdot $256$ $13.4\cdot 1024$ $17.8\cdot 1024$ $5.3\cdot 1024$
Total hrs $36$ $27$ $39$ $34.9$ $49.4$ $29.1$
CC Best $h$ 11 25 19 12 12 20
Preproc. hrs $\cdot$ CPUs 28 $\cdot$ 161 11 $\cdot$ 151 23 $\cdot$ 92 $21.5$cdot $58$ 31.6 $\cdot$ 58 23.8 $\cdot$ 64
Recover hrs $\cdot$ GPUs 0.1 $\cdot$ 256 42 $\cdot$ 256 0.9 $\cdot$ 256 $0.04\cdot 1024$ 0.1 $\cdot$ 1024 4.2 $\cdot$ 1024
Total hrs 28.1 53 34 21.5 31.7 28
@@
Parameters. The definitions of $\tau$, $B$ and $c$ given on [22, p.21] depend on the root Hermite factor $\delta_{0}$ of the short dual vectors, and a target value for $\delta_{0}$ is not provided. Hence, we make the following engineering choice, based on experiments. We observe that only short vectors with $B<Q/8$ result in MiTM searches that run in reasonable time, so we fix $B<q/8$ and compute it using the method of the [22] implementation: $B=(2+\frac{1}{\sqrt{2\pi}})\cdot\alpha q\sqrt{\frac{m}{m+n}}\cdot\frac{||\mathbf{y_{1}}||}{c}$. $\mathbf{y_{1}}$ is a short vector obtained from the scaled dual attack, and we use the average norm of all short vectors obtained from the scaled dual reduction to compute $B$. The definition of $B$ in the paper relies on $\delta_{0}$, but formulae for estimating $\delta_{0}$ are inaccurate for $\beta<50$ [20, 17]. We fix $c=10$, mirroring [22], and use $m=n$, following [22, 5].
For $\tau$ and $\zeta$, we initially use values provided by the Lattice Estimator but find experimentally that these are inaccurate. For example, $\tau=50$ short vectors are sufficient to recover secrets, compared to the hundreds estimated by Estimator. We also find that $\zeta$ values given by the estimator make the reduction of these dual lattices unreasonably slow for the small block sizes we can tractably run. We run ablation experiments across various $\zeta$ and $\beta$ values for the other two settings, and use $\zeta$ values providing the best trade off in reduction time and secret recovery. Table VIII lists our chosen $\zeta$ and $\tau$ and experimentally chosen error bound $B/q$.
Search criterion. In the [22] implementation, the correctness of a table element is assessed by computing the $L_{\inf}$ norm (e.g. largest element) of the putative short vector $\mathbf{r}^{\prime}=\mathbf{b}^{\prime}-\mathbf{A}^{\prime}s^{*}-\mathbf{A}^{\prime}s^{\dagger}$, where $\mathbf{s}^{*}$ is an element stored in the table and $\mathbf{s}^{\dagger}$ is a guess. However, we observe that $\mathbf{r}^{\prime}$ often contains outliers, so using the $L_{\inf}$ norm may yield false negatives. We instead check against the median value of $\mathbf{r}^{\prime}$, which reduces the effect of large outliers.

### v2

@@ Abstract
Lattice cryptography schemes based on the learning with errors (LWE) hardness assumption have been standardized by NIST for use as post-quantum cryptosystems, and by HomomorphicEncryption.org for performing encrypted computations on sensitive data. Thus, understanding their concrete security is critical. Most work on LWE security focuses on theoretical estimates of attack performance, which is important but may overlook attack nuances arising in real-world implementations. The sole existing concrete benchmarking effort, the Darmstadt Lattice Challenge, does not include benchmarks relevant to the standardized LWE parameter choices—such as small secret and small error distributions, and Ring-LWE (RLWE) and Module-LWE (MLWE) variants. To improve our understanding of concrete LWE security, we provide the first benchmarks for LWE secret recovery on standardized parameters, for small and low-weight (sparse) secrets. We evaluate four LWE attacks in these settings to serve as a baseline: the Search-LWE attacks uSVP [9], SALSA [51], and Cool&Cruel [44], and the Decision-LWE attack: Dual Hybrid Meet-in-the-Middle (MitM) [21]. We extend the SALSA and Cool&Cruel attacks in significant ways, and implement and scale up MitM attacks for the first time. For example, we recover hamming weight $9-11$ binomial secrets for KYBER ($\kappa=2$) parameters in $28-36$ hours with SALSA and Cool&Cruel, while we find that MitM can solve Decision-LWE instances for hamming weights up to $4$ in under an hour for Kyber parameters, while uSVP attacks do not recover any secrets after running for more than $1100$ hours. We also compare concrete performance against theoretical estimates. Finally, we open source the code to enable future research.
footnotetext: CNRS, LIP6, F-75005 Paris, France
@@ 1 Introduction
Cryptosystems based on Learning with Errors (LWE) have emerged as leading contenders. The LWE problem is: given many instances of a random vector $a$ of dimension $n$ along with a noisy inner product of $a$ with a secret vector $s$ modulo $q$, recover $s$. The hardness of LWE depends on the dimension $n$, the modulus $q$, and the distributions from which the secret and the error are drawn.
Several LWE-based cryptosystems, such as CRYSTALS-KYBER [11] were standardized by NIST, the US National Institutes of Standards and Technology [18] in 2024 for industry use in post-quantum cryptography. Furthermore, all publicly available fully homomorphic encryption (HE) libraries rely on the hardness of LWE for their security. HE was standardized by an industry consortium in 2018 [2], with recent updated analysis given in [14]. Since standardization, LWE-based cryptosystems have been incorporated into applications such as the encrypted messaging protocol, Signal [33], and LWE-based HE schemes have been used in production by Microsoft in a password-breach detection scheme [35], and in the finance industry for encrypted computations on sensitive data [31].
@@
Our Contributions. We provide concrete benchmarks for LWE attacks with parameter choices from two key real-world use cases of LWE: CRYSTALS-KYBER and Homomorphic Encryption. We implement and evaluate four concrete LWE attacks—uSVP, SALSA, Cool&Cruel, and Dual Hybrid MiTM—on these parameter settings. We then demonstrate successful secret recovery attacks on settings mimicking standardized LWE parameters, albeit with sparse secrets (see Table IX): $n=256$, $\log q=12$, Module-LWE of ranks $2$ and $3$ with binomial secret and error distributions for KYBER; and $n=1024$, $\log q=26,29$ with ternary secrets and narrow Gaussian error for HE.
We start with sparse, or low weight, secrets, to benchmark attacks which succeed on these. The challenge is to successfully recover higher Hamming weight secrets, moving towards recovery of general secrets. Prior work has demonstrated secret recovery for non-standard parameters, such as small dimension $n=100-200$ and large $\log q$, with the goal of increasing $n$ or decreasing $q$ to make the LWE problem harder. Our proposed benchmark instead fixes $n$ and $q$ as the choices standardized for KYBER and HE, and identifies Hamming weight (or secret sparsity) as the sliding hardness parameter, to be increased to reach general secrets.
Table IX shows secret recovery of hamming weight $9-11$ binomial secrets for KYBER parameters in $28-36$ hours with SALSA and Cool&Cruel, while we find that MitM can solve Decision-LWE instances for hamming weights up to $4$ in under an hour for KYBER parameters, while uSVP attacks do not recover secrets after running for more than $1100$ hours. In the process of implementing and evaluating attacks, we made many new technical contributions, including:
@@ 2.1 LWE Settings
  • Binomial, ${{\mathfrak{B}}^{\eta}}$: from a centered binomial distribution by taking independent uniform samples $(a_{1}\ldots a_{\eta},b_{1}\ldots b_{\eta})\leftarrow\{0,1\}^{2\eta}$, then outputting $\Sigma_{i=1}^{\eta}(a_{i}-b_{i})$ [11].
@@ 2.2 LWE in the real world
Homomorphic Encryption. Most publicly available HE libraries implement RLWE-based systems and often use small (binary, ternary, narrow Gaussian), sparse secrets and small error. Small, sparse secrets enable bootstrapping for evaluating deep circuits required for deep neural nets, such as in encrypted AI model inference [29]. In some proposed schemes, secrets have only $h=64$ active bits for dimension $n=2^{14}$ [22]. Table 4.2 of [14] gives the highest modulus $q$ that can be safely used for a given $n$ with ternary or narrow Gaussian secret distribution $\chi_{s}=\mathcal{N}(3.19)$ and narrow Gaussian error $\chi_{e}=\mathcal{N}(3.19)$ (see Table III).
@@ 3.1 Benchmark Settings
We propose a set of practical benchmark settings to encourage concrete evaluations of LWE attack performance, using the following criteria. First, parameters should come from real-world choices, such as narrow secret and error distributions and LWE variants like Ring- and Module-LWE, that are standardized or used in deployed LWE cryptosystems. Prior work shows that such parameter choices—including binary secrets, small error, and use of the ring-LWE variant—may weaken LWE, necessitating further experimental study [12, 28, 23, 48, 44]. The challenges should also have tunable hardness settings. The hardness of LWE problems depends on dimension $n$, modulus size $q$, error distribution $\chi_{e}$, and secret distribution $\chi_{s}$. Generally, larger dimension $n$, smaller modulus $q$, and secret/error distributions with higher standard deviation ($\sigma_{e}$, $\sigma_{s}$) make LWE more difficult. A good set of challenges should fix some parameters and allow others to vary so that concrete attacks can be bench-marked to provide interesting insights.
Using these criteria, we propose two sets of benchmark settings for measuring concrete LWE attack performance based on two important real-world applications of LWE: KYBER and Homomorphic Encryption. Security guidelines have been proposed for each of these, outlining the required LWE distribution, $n$, $q$, $\chi_{s}$, and $\chi_{e}$ for real-world implementations, as discussed in §2.2 and Tables II and III. Since $n$, $q$, $\chi_{s}$, and $\chi_{e}$ are fixed for KYBER and for $n=1024$, $\chi_{s}$ and $\chi_{e}$ are fixed for HE, our challenges vary problem difficulty by changing the number of nonzero secret coordinates, $h$, and modulus, $q$. We divide our proposed benchmark settings into parameters focused on KYBER, using MLWE, and parameters focused on HE, using RLWE, and present the proposed settings in Tables VI and VI.
The challenge is to solve LWE problems for larger $h$ in these standard parameter settings. Prior work has extensively explored attack performance for small $n$ with uniform secrets (e.g.  [15]). In addition, it is already well-known that lattice problems such as LWE are not hard when $\log q$ becomes large enough, for any fixed $n$ [34, 23]. Our challenge instead encourages implementation of full-scale attacks on large $n$ and small $q$, for small secret distributions which are standardized but potentially more risky. We tune hardness by increasing the number of non-zero elements in the secret.
@@ 3.2 Choosing Attacks to Evaluate
To choose which attacks to evaluate, we focused on: (1) attacks which have open-source implementations, (2) attacks which have already been evaluated in nontrivial settings (e.g. dimension $n>256$), (3) attacks that require less than $750$ GB of RAM (our machine capacity). We chose to evaluate the 4 attacks discussed in Section 4: uSVP, ML, and Cool&Cruel for Search-LWE and Dual Hybrid MitM for Decision-LWE. Our attack evaluation provides a first set of benchmarks for the proposed settings. Future work should improve these and evaluate additional attacks.
@@ 4 Attack Implementations and Innovations
Here, we present details of the attacks we evaluate, as well as the innovations we introduce to make these attacks run on the proposed benchmark settings. Refer to the original attack papers for details. All attacks are implemented in a codebase available at https://github.com/facebookresearch/LWE-benchmarking. Implementations are written in Python and leverage fplll and flatter [49] libraries for lattice reduction (with enumeration as the SVP oracle in BKZ2.0). Since all attacks rely on lattice reduction, benchmarking them using the same lattice reduction implementation allows for fair comparison. Any improvement to lattice reduction would benefit all attacks.
@@
The space spanned by these rows contains an unusually short vector $\begin{pmatrix}s&-e&-1\end{pmatrix}$.
Thus lattice reduction recovers the secret once it finds the shortest vector in the lattice.
Implementation details. Our implementation of uSVP multiplies the $I_{n}$ matrix by a factor $\omega$, which balances $s$ and $-e$ in the discovered short vector. $\omega$ is determined by formulae given in [23]. We run lattice reduction using BKZ2.0 [20] and incorporate the improvements suggested in [38, Appendix A.7, p. 17]. Future improvements to our implementation of the uSVP attack could incorporate more advanced BKZ schemes, such as Pump and Jump [52].
@@ 4.2 ML Attack
However, we observe experimentally that since $\delta\bf{e_{i}}$ falls outside of $f$’s training distribution, $f$ does not produce helpful predictions on this input. Thus, we propose an alternative secret recovery method that embodies this concept, while remaining within the data distribution. We call it a “slope distinguisher.” This distinguisher calculates the slopes, or approximations of the derivatives, of model outputs using the formula $\frac{\partial f}{\partial x_{i}}(a)\approx\frac{f(a+\delta e_{i})-f(a)}{\delta}=:\hat{s}_{i}$ for some $a$ drawn from the data distribution. To account for the noisy predictions, we compute many samples of $\hat{s}_{i}$ and take the most frequently appearing value, rounded to the nearest integer. With this, the ML method can recover general secrets. This new distinguisher is illustrated in Figure 1.
@@ 4.4 Dual Hybrid MiTM
ML Best $h$ 9 18 16 10 10 17
Preproc. hrs $\cdot$ CPUs $28$cdot $3216$ $11$cdot $3010$ $33$cdot $1843$ $32$cdot $1160$ $31.6$cdot $1164$ $23.8$cdot $1284$
Recover hrs $\cdot$ GPUs $8$cdot $256$ $16$cdot $256$ $6$cdot $256$ $14\cdot 1024$ $17.8\cdot 1024$ $5.3\cdot 1024$
Total hrs $36$ $27$ $39$ $46$ $49.4$ $29.1$
CC Best $h$ 11 25 19 11 12 20
Preproc. hrs $\cdot$ CPUs 28 $\cdot$ 161 11 $\cdot$ 151 23 $\cdot$ 92 $32$cdot $58$ 31.6 $\cdot$ 58 23.8 $\cdot$ 64
Recover hrs $\cdot$ GPUs 0.1 $\cdot$ 256 42 $\cdot$ 256 0.9 $\cdot$ 256 $0.1\cdot 1024$ 0.1 $\cdot$ 1024 4.2 $\cdot$ 1024
Total hrs 28.1 53 34 32.1 31.7 28
@@
Parameters. The definitions of $\tau$, $B$ and $c$ given on [21, p.21] depend on the root Hermite factor $\delta_{0}$ of the short dual vectors, and a target value for $\delta_{0}$ is not provided. Hence, we make the following engineering choice, based on experiments. We observe that only short vectors with $B<Q/8$ result in MiTM searches that run in reasonable time, so we fix $B<q/8$ and compute it using the method of the [21] implementation: $B=(2+\frac{1}{\sqrt{2\pi}})\cdot\alpha q\sqrt{\frac{m}{m+n}}\cdot\frac{||\mathbf{y_{1}}||}{c}$. $\mathbf{y_{1}}$ is a short vector obtained from the scaled dual attack, and we use the average norm of all short vectors obtained from the scaled dual reduction to compute $B$. The definition of $B$ in the paper relies on $\delta_{0}$, but formulae for estimating $\delta_{0}$ are inaccurate for $\beta<50$ [19, 23]. We fix $c=10$, mirroring [21], and use $m=n$, following [21, 5].
For $\tau$ and $\zeta$, we initially use values provided by the Lattice Estimator but find experimentally that these are inaccurate. For example, $\tau=50$ short vectors are sufficient to recover secrets, compared to the hundreds estimated by Estimator. We also find that $\zeta$ values given by the estimator make the reduction of these dual lattices unreasonably slow for the small block sizes we can tractably run. We run ablation experiments across various $\zeta$ and $\beta$ values for the other two settings, and use $\zeta$ values providing the best trade off in reduction time and secret recovery. Table VIII lists our chosen $\zeta$ and $\tau$ and experimentally chosen error bound $B/q$. Search criterion. In the [21] implementation, the correctness of a table element is assessed by computing the $L_{\inf}$ norm (e.g. largest element) of the putative short vector $\mathbf{r}^{\prime}=\mathbf{b}^{\prime}-\mathbf{A}^{\prime}s^{*}-\mathbf{A}^{\prime}s^{\dagger}$, where $\mathbf{s}^{*}$ is an element stored in the table and $\mathbf{s}^{\dagger}$ is a guess. However, we observe that $\mathbf{r}^{\prime}$ often contains outliers, so using the $L_{\inf}$ norm may yield false negatives. We instead check against the median value of $\mathbf{r}^{\prime}$, which reduces the effect of large outliers.